# Word length predicts word order: “Min-max”-ing drives language evolution

### v2

@@ 3.1 A “Min-Max” theory of processing
Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate. In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions. Similarly, while all SV languages have a high proportion of N1 sentences (max 2.43 n/v, 7.17 a/p), the number of V1 constructions is non-zero, supporting a gradient approach to word order (Levshina et al. 2023).
The fact that the length of particular word classes can predict the basic word order of a language is a novel finding, leading to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order. If, on the other hand, verbs are longer, it is highly likely to have VS word order. The effect of frequency cannot be ignored either, highlighting the importance of conducting large-scale corpus based studies to investigate this question. Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically. The pressures identified in this paper have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
@@ 3.2 Limitations and future directions
One limitation in this set of studies is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.

### v3

@@ 3.1 A “Min-Max” theory of processing
The importance of frequency to any account of language evolution is that frequency reflects usage. The most frequent items in a corpus are those that are most used by speakers of a language. As such, under a statistical learning paradigm, they are most likely to influence generalizations made by speakers/hearers based on their input. This means that any discussion of linguistic forms over time ought to consider frequency as a factor in how such forms change.
As an example to illustrate the potential impact of this theory, I offer the following thought experiment involving a hypothetical language. This hypothetical language is limited to nouns and verbs, which the speakers use in their interactions, in goal-directed behavior such as making tools. Nouns and verbs are the same length, and the informational content of each word is simply based on their communicative impact in the interactions. As speakers use the nouns and verbs to communicate about their tool-making, there is no pressure from noun/verb lengths on order of constituents, with order being determined by the need to identify objects/actions in immediate interaction. Over time, as speakers innovate and additional information begins to get encoded (such as deixis), the information will become associated with either nouns or verbs, possibly resulting in additional morphology, and potential affixation. The relevant noun/verb associated with such morphology then becomes “heavier” and provides more information for the interaction. This makes them more likely to be placed at the beginning of an utterance, since that makes it easier to predict what follows. The information-rich noun/verb has great utility and therefore has high frequency, becoming notable and relevant to statistical inference by language learners, who may then generalize an ordering rule that places the heavier constituent at the front of the speech stream.
How does this relate to language change? In this thought experiment, we could consider what might happen if a language that had developed markers on verbs only (and therefore had VS order) innovated a new set of tools for particular tasks, but with only a single word for “tool”. As the tools became more prevalent, usage increased in interaction, and each tool began to acquire affixes that identified their specific use and other relations. This developed to the point that a listener could infer, from the named object, a particular context and subsequent utterances. At this point the nouns have become “heavier” than the verbs, with the resulting stronger pressure to place them before verbs in the speech stream, and the language would realize SV order.
While this is only a thought experiment, it does align with what we know about the development of pidgins and creoles. Pidgins are languages that arise out of contact situations and tend to lack morphology. These can develop into creoles as they acquire native speakers (children who grow up speaking the pidgin), and they begin to exhibit increased morphological complexity as the younger speakers innovate forms. The possibility laid out in this thought experiment is also supported by the finding in this paper that the length of particular word classes can predict the basic word order of a language. In turn, this leads to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order. If, on the other hand, verbs are longer, it is highly likely to have VS word order. The effect of frequency highlights the importance of conducting large-scale corpus based studies to investigate this and related concerns. Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
@@
Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate. In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions, and suggesting that there is a strong efficiency pressure toward SV order. Similarly, while all SV languages have a high proportion of N1 sentences (max 2.43 n/v, 7.17 a/p), the number of V1 constructions is non-zero, supporting a gradient approach to word order (Levshina et al. 2023).
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which may then lead to different realizations of word order crosslinguistically. The pressures identified in this paper have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
@@ 3.2 Limitations and future directions
One limitation in this set of studies is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2026 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.